OpenStudy (anonymous):

Simplify (3x^0y^7)(2x^2y)^3

OpenStudy (blurbendy):

start by cubing everything on the left side what is 2^3?

OpenStudy (blurbendy):

i mean right side

OpenStudy (anonymous):

2^3=8

OpenStudy (blurbendy):

yes, (x^2)^3 ?

OpenStudy (anonymous):

that is x^6

OpenStudy (blurbendy):

yes and (y)^3 ?

OpenStudy (anonymous):

thats y^3

OpenStudy (anonymous):

\[ (3x^0y^7)(2x^2y)^3= (3x^0y^7)(2^3x^{(2\times 3)}y^3)=(3x^0y^7)(8x^{6}y^3)\] \[=3\times 8x^{0+6}y^{7+3} =24x^{6}y^{10}\] is the required simplified form

OpenStudy (blurbendy):

yes, so you have (3x^0y^7)(8x^6y^3) now, just multiply

OpenStudy (anonymous):

thanks both of you

OpenStudy (blurbendy):

yw